OpenStudy (anonymous):

solve each equation on the interval [0,2pi] cosx-2sinxcosx=0

OpenStudy (anonymous):

\[\begin{align*} \cos x-2\sin x\cos x&=0\\ \cos x(1-2\sin x)&=0\end{align*}\\ \implies\begin{cases}\cos x=0\\1-2\sin x=0\end{cases}\]

OpenStudy (anonymous):

hello?